OpenStudy (anonymous):

At a large nursery, a border for a rectangular garden is being built. Designers want the border's length to be 5 ft greater than it's width. A maximum of 180 ft of fencing is available for the border. Write and solve an inequality that describes possible widths of the garden.

OpenStudy (anonymous):

please help!!!!!!!!!!!!!

OpenStudy (anonymous):

Perimeter of a rectangle = 2w+2l Perimeter = 180ft Width = w Length=w+5 Can you write an equation from this?

OpenStudy (anonymous):

Write an equation that involves the perimeter of the rectangle. Use the values and expressions above.

OpenStudy (anonymous):

\[2(2w+5)\leq 180 \]\[w\leq \frac{85}{2} \]

OpenStudy (anonymous):

What is the point if you just giving the answer, without walking the person thru it so they can work thru the problem.

OpenStudy (anonymous):

Sabrinaaa r u still there?

OpenStudy (anonymous):

yeah sorry just trying to think

OpenStudy (anonymous):

w+5 > 180 with the underline on the equality sign? yes? or no?

OpenStudy (anonymous):

Were you able to find the equation for the perimeter of the rectangle? Perimeter <=180ft Width=w Length=w+5 Using the Formula to find the perimeter of a rectangle: P=2l+2w, you should get: 2(w+5)+2w<=180 Do you understand this?

OpenStudy (anonymous):

w=42.5

OpenStudy (anonymous):

Yes to a certain extent. The actual answer is: w<=42.5

OpenStudy (anonymous):

wait what?

OpenStudy (anonymous):

Remember it is an inequality question. The maximum perimeter of the rectangle is 180ft. from the inequality equation:2(w+5)+2w<=180, when you solve for w you should get w<=42.5 not an exact value.

OpenStudy (anonymous):

ok i kinda get it now

OpenStudy (anonymous):

Good. So the value of the width can be 42.5 or a number less than that.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

sure no problem